# Bridging the Gap: Towards an Expanded Toolkit for ML-Supported Decision-Making in the Public Sector

### v2

@@ 3.3.1 Distribution Shift, Selection Bias and Label Bias
As described, a significant challenge in managing distribution shifts for ADM systems lies in precisely specifying the anticipated changes from the historical environment to the future deployment environment. For example, the data-generating mechanism may change over time, but the exact nature of this shift is usually hard to predict. Addressing this challenge, Si et al., (2020, 2023) propose an algorithm for distributionally robust policy learning under unknown covariate and concept shift. Their approach involves maximizing the worst-case policy value over all environments within a specific distance to the training environment. Optimizing over a large neighborhood of distributions guarantees robustness against significant shifts, yet may substantially reduce the performance of the policy if the environment changes less than anticipated. By choosing their preferred distance, decision-makers can manage their risk aversion before deploying a policy (Si et al., , 2023). Building on this work, Kallus et al., (2022) incorporate DR methods, removing the need to assume that the historical assignment policy is known, which is often unavailable when relying on observational data.
However, domain experts may often encounter difficulties in precisely specifying the optimal neighborhood size (Hatt et al., , 2022). Instead of ensuring robustness under arbitrary distribution shifts, it may be helpful to focus on specific types of shifts, potentially simplifying the integration of domain knowledge. For example, Hatt et al., (2022) develop a framework for learning worst-case policies that generalize under distributional shifts resulting from an unknown selection bias.
@@ 3.3.2 Uncertainty Quantification
After selecting a policy for deployment, especially in high-stakes settings, reliable uncertainty estimates become important to guarantee the policy’s reliability. Uncertainty quantification in off-policy learning often involves estimating bounds for the expected aggregate utility of the policy under hypothetical deployment (Taufiq et al., , 2022), for example as seen in Wang et al., (2017). However, in many scenarios there may arise the need to quantify the uncertainty of outcomes at the individual level. For instance, a policy that appears to lead to a positive aggregate utility may still be deemed unacceptable if the variability in outcomes for certain subgroups is overly large. Recent works in off-policy evaluation have investigated the application of conformal prediction to construct prediction intervals for individual covariates. Similar to CATE estimation, a critical challenge for conformal off-policy prediction lies in guaranteeing exchangeability of the data. Zhang et al., (2023) and Taufiq et al., (2022) have proposed approaches that make use of weighted conformal prediction to address the shift between training data and deployment environment, allowing for the reliable estimation of prediction sets.
@@ 3.3.3 Multi-Objective Optimization
While there is significant research on multi-objective reinforcement learning (Hayes et al., , 2022), the focus often lies on online policy learning and less attention is given to multi-objective policy learning from historical data. To generalize the policy value for multiple objectives, one can consider the weighted sum of utilities resulting from various individual outcomes and external objectives. For example, in scenarios with individually varying intervention costs, it may be possible to define a net-monetary benefit, that is subsequently used as the target outcome in the policy value (Xu et al., , 2022). Alternatively, a decision-maker may want to enforce an overarching constraint, such a limited budget, as part of the policy optimization problem (Huang and Xu, , 2020; Sun, , 2021). However, if the relevant constraint needs to be adjusted regularly, such approaches may lead to significant computational costs (Sun et al., , 2024). Sun et al., (2024) propose learning an individual-level priority score that directly encodes the cost-benefit ratio, which can subsequently easily be used to rank individuals for intervention under varying resource constraints.
Off-policy learning methods that optimize within a constrained class of policies (Athey and Wager, , 2021; Kallus, , 2021) have the advantage that secondary constraints can also be encoded through restricting the class of allowed policies. For example, Frauen et al., (2023) propose a policy learning that restricts the policy class to those respecting fairness constraints.
@@ 4 Discussion
Making productive use of ML for public sector decision-making is a complicated task, requiring careful alignment of policy objectives and model development. We provided a overview of technical challenges that commonly arise when aligning policy objectives with ML models with the goal of enhancing decision-making in the public sector. We discussed the intricacies of acquiring representative training data and addressing issues like robustness under distribution shifts, dealing with proxy variables, and the influence of past decision-making. We also covered various challenges complicating the link between policy goal, model output and decision-making, including the difficulties of addressing multiple objectives and human involvement in decision-making.
Furthermore, we examined key target estimands and their alignment with decision-making scenarios in the public sector: risk prediction, counterfactual prediction and policy learning. These modeling frameworks do not only add structure to decide and evaluate on how a given policy problem may be best connected or supported by ML approaches, but also point to context-specific solutions to the practical challenges outlined above. We highlighted relevant methodological advancements for causal estimation, handling distribution shifts, uncertainty quantification and multi-objective optimization that extend beyond the scope of the standard ML framework. We provide a distilled version of our presentation of different target estimands, modeling approaches and their links to policy objectives and decision-making in Table 1, with the aim of providing guidance for discussions on which modeling theme is most suitable in a given scenario. We recommend that policymakers and model developers collaboratively select the modeling approach most suited for their intended application by establishing a clear connection between the policy goal, allocation principle and ultimately the target estimand.
@@
The intention behind this paper was to work toward the development of a robust methodological framework for constructing and maintaining ADM systems in the public sector that includes both best practices for practitioners and an up-to-date array of technical approaches. While we have made some inroads here by selecting and consolidating relevant theoretical advancements, a critical need for more research to connect these methods with real-world policy use cases remains. As illustrated by the methods and challenges presented here, achieving this goal will likely require bringing together researchers from various disciplines to develop systems that genuinely improve decision-making in tangible ways. It will require a shift in perspective away from technical approaches purely centered around predictive optimization and towards ones that explicitly incorporate decision-making and its impact into the modeling framework.
@@ Acknowledgment
This work is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research, by the Baden-Württemberg Foundation under the grant “Fairness in Automated Decision-Making - FairADM” and by the Volkswagen Foundation, grant “Consequences of Artificial Intelligence for Urban Societies (CAIUS)”. We express our gratitude to Felix Henninger, Patrick Schenk, Nanina Föhr and Daria Szafran for their valuable feedback and support, and to Bernd Fitzenberger, for shaping our thinking about applications of prediction models in the public sector.
Appendix A CATE Estimation Methods
@@
Alternative methods directly estimate the CATE function by first constructing pseudo-outcomes of the treatment effects (Künzel et al., , 2019; Caron et al., , 2022; Curth and van der Schaar, , 2021). One prominent variant is the X-learner (Künzel et al., , 2019), an extension of the T-learner, which can also be regarded as a special case of the RA-learner (Curth and van der Schaar, , 2021). The Doubly-Robust learner (DR-learner) employs pseudo-outcomes constructed from both the conditional outcomes $\hat{\mu}_{t}(x)$ and the propensity scores $\hat{\pi}(x)$ (Kennedy, 2020b, ). DR estimators have a long history in causal inference and missing data imputation (Kang and Schafer, , 2007; Bang and Robins, , 2005; Robins et al., , 1994; Funk et al., , 2011) and offer the advantage of consistency as long as either the propensity score model or conditional outcome models are correctly specified. Finally, the R-learner (Nie and Wager, , 2020; Foster and Syrgkanis, , 2023) involves the formulation a specific loss function after fitting several nuisance functions that can be separately minimized and regularized to estimate the CATE, drawing inspiration from the Robinson decomposition (Robinson, , 1988). There also exists CATE estimation methods that adapt specific ML models (Jacob, , 2021). For instance, causal forests, introduced by Athey et al., 2019b ; Wager and Athey, (2018), resemble the R-learner (Caron et al., , 2022; Post et al., , 2022; Oprescu et al., , 2019).
@@
Appendix B Off-Policy Learning

### v3

@@ 3.3.1 Distribution Shift, Selection Bias and Label Bias
As described, a significant challenge in managing distribution shifts for ADM systems lies in precisely specifying the anticipated changes from the historical environment to the future deployment environment. For example, the data-generating mechanism may change over time, but the exact nature of this shift is usually hard to predict. Addressing this challenge, Si et al., (2020, 2023) propose an algorithm for distributionally robust policy learning under unknown covariate and concept shift. Their approach involves maximizing the worst-case policy value over all environments within a specific distance to the training environment. By choosing their preferred distance, decision-makers can manage their risk aversion before deploying a policy (Si et al., , 2023). Building on this work, Kallus et al., (2022) incorporate doubly-robust methods, removing the need to assume that the historical assignment policy is known, which is often unavailable when relying on observational data. Instead of ensuring robustness under arbitrary distribution shifts, it may also be helpful to focus on specific types of shifts, potentially simplifying the integration of domain knowledge. For example, Hatt et al., (2022) develop a framework for learning worst-case policies that generalize under distributional shifts resulting from an unknown selection bias.
@@ 3.3.2 Uncertainty Quantification
After selecting a policy for deployment, especially in high-stakes settings, reliable uncertainty estimates become important to guarantee the policy’s reliability. Uncertainty quantification in off-policy learning often involves estimating bounds for the expected aggregate utility of the policy under hypothetical deployment (Taufiq et al., , 2022), for example as seen in Wang et al., (2017). However, in many scenarios there may arise the need to quantify the uncertainty of outcomes at the individual level. For instance, a policy that appears to lead to a positive aggregate utility may still be deemed unacceptable if the variability in outcomes for certain subgroups is overly large. Recent works in off-policy evaluation have investigated the application of conformal prediction to construct prediction intervals. Similar to CATE estimation, a critical challenge for conformal off-policy prediction lies in guaranteeing exchangeability of the data. Zhang et al., (2023) and Taufiq et al., (2022) have proposed approaches that make use of weighted conformal prediction to address the shift between training data and deployment environment, allowing for the reliable estimation of prediction sets.
@@ 3.3.3 Multi-Objective Optimization
To generalize the policy value for multiple objectives, one can consider the weighted sum of utilities resulting from various individual outcomes and external objectives. For example, in scenarios with individually varying intervention costs, it may be possible to define a net-monetary benefit, that is subsequently used as the target outcome in the policy value (Xu et al., , 2022). Alternatively, a decision-maker may want to enforce an overarching constraint, such a limited budget, as part of the policy optimization problem (Huang and Xu, , 2020; Sun, , 2021). However, if the relevant constraint needs to be adjusted regularly, such approaches may lead to significant computational costs (Sun et al., , 2024). Sun et al., (2024) propose learning an individual-level priority score that directly encodes the cost-benefit ratio, which can subsequently easily be used to rank individuals for intervention under varying resource constraints. Furthermore, off-policy learning methods that optimize within a constrained class of policies (Athey and Wager, , 2021; Kallus, , 2021) have the advantage that secondary constraints can also be encoded through restricting the class of allowed policies. For example, Frauen et al., (2024) propose a policy learning that restricts the policy class to those respecting fairness constraints.
@@
4 Towards a Decision-Centric ML Toolkit in the Public Sector
Making productive use of ML for public sector decision-making is a complicated task, requiring careful alignment of policy objectives and model development. First, we set out to explore challenges faced when deploying ML for public sector decision-making. Specifically, we focused on challenges arising from a misalignment between policy objectives and technical design, such as when assumptions about the training data do not match the intended application context. We identified and discussed five key challenges in Section 2, highlighting potential limitations of solely relying on the standard supervised ML paradigm (see Figure 2). In response to these limitations, we examined alternative modeling frameworks, specifically counterfactual prediction and off-policy learning. Each framework comes with its own set of distinct advantages and is potentially better suited to overcome some of the challenges. To choose between these frameworks, a model developer should consider two key questions. 1) How will the estimated quantity be helpful in decision-making? 2) What is necessary to ensure unbiased estimation?
We observe that targets that are easier to estimate, often require more assumptions about their utility in the decision-making process. Even with perfect knowledge of these targets, they might need to be combined with domain knowledge to be informative for the decision-maker. Conversely, estimating causal outcomes requires more assumptions for causal identification, but might offer more actionable insight for decision-makers. We provide a distilled version of our presentation of different modeling approaches and their links to policy objectives and decision-making in Table 1, with the aim of providing guidance for discussions on which modeling theme is most suitable in a given scenario.
@@ 4 Towards a Decision-Centric ML Toolkit in the Public Sector
For example, traditional (risk) prediction methods, while requiring fewer assumption during the estimation process, are not well-suited to estimate counterfactual outcomes, which are often the true target of interest for decision-makers. This limitation often requires additional assumptions about how the predictions relate to the decision-maker’s objective. Consider the previously discussed medical scenario as an example for such an assumption: the decision-maker assumes that a higher mortality risk in the coming years might make a knee replacement less beneficial, while also assuming that the mortality risk is not significantly impacted by the surgery itself. Counterfactual prediction can potentially support a variety of decision-makers goals with fewer explicit assumptions about how the goals of the decision-maker align with the target of estimation. For example, consider a public employment service aiming to target support measures to job seekers with a high risk of becoming long-term unemployed. Similarly, reliable estimates of the heterogeneous causal effects of a support program would aid a decision-maker in matching individual’s to the most effective program. However, counterfactual estimation is more involved than standard (risk) prediction, relying on external assumptions to ensure causal identification. Policy learning is explicitly integrated into the decision-making process by optimizing a predefined utility function to directly estimate an allocation policy. Such an end-to-end approach allows for a more straightforward integration of constraints and additional objectives. However, it might struggle in scenarios in which human decision-makers play a crucial role. In such settings, decision-makers might prefer individual-level estimates as recommendations to guide their own judgments. We present three examples inspired by real-world use cases in Figures 3, 4 and 5, each focusing on risk prediction, counterfactual prediction and policy learning respectively. They summarize some of the key questions practitioners need to address to ensure that the selected approach fits the intended application context.
Certain challenges such as the influence of past decision-making are inherently addressed by causal modeling frameworks. To tackle the remaining challenges, we have compiled for each modeling approach a selection of methods to address them, as detailed in the previous section and summarized in Table 3 in the appendix. Our goal was to identify methods that are applicable across various ML models within each respective modeling framework, to keep most of our discussion model agnostic and relevant across many application contexts. Unsurprisingly, there is generally less research addressing specific challenges within newer causal modeling frameworks. This gap presents a compelling direction for future research to explore which mitigation strategies from standard supervised ML could be extended to the causal setting.
In recent years, several studies have applied causal ML frameworks to practical applications in the public sector. For example, these modeling approaches have been used and evaluated for the optimal allocation of development aid (Kuzmanovic et al., , 2024), allocation of medical preventive care (Kraus et al., , 2024), child welfare hotline screening (Coston et al., , 2020), and assignment of training programs to job seekers (Cockx et al., , 2023).
@@ 5 Discussion
We analyzed challenges that result from a misalignment between the (technical) assumptions made during model design and the intended policy goals. Each challenge can lead to harmful unintended consequences, impacting the individuals affected by the decisions and potentially undermining the legitimacy of the system.
  • Distribution Shifts occur when the data used to train the ML model does not reflect real-world conditions, causing the performance of the model to decline. Such shifts can lead to misclassifications of individuals and create harmful feedback loops that reinforce erroneous predictions.
  • Label Bias can happen when a ML model relies on proxy variables to estimate hard to measure outcomes. If these proxies are biased and primarily reflect institutional practices instead of of the true target, they can systematically disadvantage certain groups, leading to unfair predictions and decision outcomes.
  • Past Decision-Making often influences the available training data. If we do not explicitly account for these past interventions, the predictions of the ADM system may become outdated once new decision-making practices are implemented. For example, a model might underestimate the risk for individuals who previously received support, potentially leading to harm if future allocations fail to consider this.
  • Competing Objectives and Constraints can complicate the formalization of policy goals in an ADM system, as predictive systems often focus on singular, clearly defined objectives. This narrow focus can introduce omitted-payoff bias, such as when optimizing for cost-efficiency disproportionately harms marginalized groups.
  • Human-in-the-Loop is an important component of ADM systems, because automated systems alone often do not meet all necessary requirements for real-world deployment. However, interactions between models and human decision-maker can introduce complications and biases. For example, an accurate prediction may not be helpful if case workers lack trust due to unclear communication of model uncertainties.
In this work, we found that standard machine learning approaches are not necessarily well-suited for the public sector context, requiring model designers to expand their toolkit to effectively address these issues. We discussed different technical solution strategies in detail and provided guidance on choosing between alternative modeling frameworks - including predictive and causal modeling - to tackle these challenges. However, to do so effectively, the technical model design needs to be guided in close collaboration with domain experts. Each challenge we presented is embedded in complex, changing social contexts, where purely technical solutions often fall short. The right technical design choice often has no clear or definite answer and can not be left to the model developer alone. For example, the implicit assumptions a model developer makes about the data distribution, the causal structure of the problem, or how different objectives should be represented in the system can greatly affect the validity of the resulting system. However, these assumptions often need to be informed domain-specific knowledge, necessitating insights from policy makers, social scientists and other stakeholders involved. In Table 2, we summarize technical solution strategies for each challenge and highlight the points where external input may be central to ensure that technical design choices remain aligned with real-world policy objectives. However, engaging and collaborating with stakeholders is often difficult in practice. Participatory design of AI systems is often mentioned as an important approach, but can be difficult to implement effectively. By specifying the points along the ML pipeline where stakeholders collaboration is crucial for supporting technical design decisions, we hope to guide these efforts and make participatory approaches more actionable (Delgado et al., , 2023).
Challenge Technical Solution Strategies Stakeholder Input
Distribution Shifts Apply domain adaptation methods using data from the deployment environment Use distributionally robust optimization for worst-case guarantees Implement continuous monitoring of input data and model performance Collaborate with domain experts and data providers to anticipate changes in the deployment environment (e.g. changing regulations and user behavior) Engage stakeholders to determine acceptable risk tolerance Involve decision-makers to determine relevant evaluation metrics for ongoing monitoring Identify vulnerable and hard-to-reach sub-populations
Label Bias Construct measurement error models for selected proxy variables Validate chosen proxy variables using external data sources and additional variables Collaborate with domain experts to understand how proxy variables map to the true concepts of interest Identify societal biases that may impact the proxy-target relationship
Past Decision-Making Identify suitable (causal) estimands and estimation strategies, such as CATE estimation, counterfactual prediction or policy learning Validate link that the chosen (causal) estimand is able to inform the decision-making process Make use of domain expertise to inform assumptions necessary for causal identification, such as gathering knowledge on past decision-making criteria and processes
Competing Objectives and Constraints Integrate external constraints into model design and allocation procedure (e.g. using model multiplicty and constrained optimization) Identify solutions along the Pareto frontier to enable decision-makers to manage tradeoffs Elicit preferences from decision-makers to quantify tradeoffs between different objectives and constraints Collaborate with stakeholders to identify objectives not fully captured by the ADM system
Human-in-the-Loop Use uncertainty quantification (e.g. conformal prediction) and explainable ML methods to provide guarantees to decision-makers and enhance transparency Understand how model outputs will be interpreted and used by decision-makers, considering user background and workflows Regularly gather feedback from end-users to improve model integration
Table 2: Overview of key challenges in ADM systems and technical solution strategies, focusing on the role of domain expertise in guiding technical design choices.
However, our focus does not cover all potential issues related to the use of predictive algorithms in the public sector. While ensuring that a system accurately reflects the goals of decision-makers is a prerequisite for ethical and reliable use, this alone is not sufficient. We do not explicitly discuss ethical, legal and broader societal challenges. Even if a system functions perfectly according to its intended goals, it can still lead to adverse outcomes. For example, this can happen if a decision-maker does not prioritize fair treatment of sensitive subgroups as an explicit design goal (Barocas et al., , 2023).
@@
The intention behind this paper was to clarify the assumptions behind different (predictive) modeling approaches, and help practitioners identify where common technical assumptions may not hold true in the public sector context. We see this as a first step towards the development of a robust methodological framework for constructing and maintaining ADM systems in the public sector that includes both best practices for practitioners and an up-to-date array of technical approaches. While we have made some inroads here by selecting and consolidating relevant theoretical advancements, a critical need for more research to connect these methods with real-world policy use cases remains. As illustrated by the methods and challenges presented here, achieving this goal will likely require bringing together researchers from various disciplines to develop systems that genuinely improve decision-making in tangible ways. It will require a shift in perspective away from technical approaches purely centered around predictive optimization and towards ones that explicitly incorporate decision-making and its impact into the modeling framework.
@@
6 Conclusion
In this paper, we analyzed misalignments between the assumptions underlying ML models and the realities of public sector decision-making. We isolated and discussed five central challenges: distribution shifts, label bias, the influence of past decision-making, competing objectives and constraints and the integration of human decision-makers. We demonstrated how misalignment can lead to unreliable and harmful predictions, potentially causing systems to fail in achieving the intended policy goals and undermining the legitimacy of the decision-making process. Through our analysis, we concluded that many assumptions commonly made in the implementation of ML models do not hold in complex, evolving decision-making environments. In response, we argue for a shift in modeling efforts from focusing solely on predictive accuracy to improving decision-making outcomes. We presented alternative modeling approaches, including causal machine learning methods including counterfactual prediction and policy learning, which may be better suited to inform decision-making. We also provided guidance on selecting the appropriate modeling strategy by clarifying the assumptions underlying these approaches. Model developers should carefully consider how the estimated quantities can guide decision-making and whether unbiased estimation is possible given available data and external assumptions. Additionally, we summarized technical solutions to the discussed challenges, such as distributionally robust optimization, uncertainty quantification and multi-objective optimization within each modeling framework. Finally, we found that selecting the right methods and frameworks requires external input from domain experts and stakeholders to ensure that the implicit assumptions made by model developers align with the specific problem setting.
@@ Acknowledgment
This work is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial In- telligence, sponsored by the Federal Ministry of Education and Research, by the Baden-Württemberg Foundation under the grant “Fairness in Automated Decision-Making - FairADM” and by the Volkswagen Foundation, grant “Consequences of Artificial Intelligence for Urban Societies (CAIUS)”. We express our gratitude to Nanina Föhr, Felix Henninger, Patrick Schenk, and Daria Szafran for their valuable feedback and support, and to Bernd Fitzenberger, for shaping our thinking about applications of prediction models in the public sector.
Appendix A Methodological Approaches for ADM Systems in Public Sector Decision-Making
Risk Prediction Counterfactual Prediction Off-Policy Learning
Distribution Shifts Biases in Data Collection (Groves and Lyberg, , 2010; Biemer, , 2010; Cornesse et al., , 2020; Yang and Kim, , 2020; Elliott and Valliant, , 2017; Kim et al., , 2022) Domain Adaptation Using Data from Deployment Environment (Kouw and Loog, , 2018; Shimodaira, , 2000; Fang et al., , 2020) Worst-Case Guarantees (Duchi and Namkoong, , 2021; Wen et al., , 2014; Zhang et al., , 2021) Shifts induced by Model Predictions (Pagan et al., , 2023; Perdomo et al., , 2020; Kim and Perdomo, , 2023) Covariate Shift between Intervention Groups (Johansson et al., , 2022; Shalit et al., , 2017; Assaad et al., , 2021) Worst-Case Guarantees (Kern et al., , 2024; Jeong and Namkoong, , 2020) Worst-Case Guarantees (Si et al., , 2023, 2020; Kallus et al., , 2022; Hatt et al., , 2022)
Label Bias Class-Conditional Label Noise (Natarajan et al., , 2013; Yu et al., , 2020) Feature-Conditional Label Noise (Wang et al., , 2021; Chen et al., , 2021) Multiple Proxy Variables (Boeschoten et al., , 2021) Counterfactual Prediction under Measurement Error (Guerdan et al., 2023a, ; Guerdan et al., 2023b, ) No dedicated methods specific to policy learning have been identified. Methods from other frameworks may be applicable.
Past Decision-Making Unbiased estimation not possible if the outcome was influenced by interventions. Requires causal identification. See Appendix B for an overview of ML-based CATE estimation methods. Requires causal identification. See Appendix C for an overview of off-policy learning methods.
Competing Objectives & Constraints Scalar Utility Functions (Keeney and Raiffa, , 1993; Boutilier, , 2013) Solutions along the Pareto Frontier (Hertweck et al., , 2022; Pfisterer et al., , 2019) and Model Multiplicity (Watson-Daniels et al., , 2023) Specific Model Constraints, such as Fairness (Hort et al., , 2023; Hardt et al., 2016b, ) and Interpretability (Molnar, , 2022) Budget-Constrained Allocation (Tu et al., , 2021; Ai et al., , 2022; McFowland III et al., , 2021) Fairness Constraints (Kim and Zubizarreta, , 2023; Mishler et al., , 2021) Budget-Constrained Allocation (Xu et al., , 2022; Huang and Xu, , 2020; Sun, , 2021) Solutions along the Pareto Frontier (Rehill and Biddle, , 2022) Specific Model Constraints (Athey and Wager, , 2021), such as Fairness (Frauen et al., , 2024)
Uncertainty Estimation for Human-in-the-Loop Model agnostic Uncertainty Estimation with Conformal Prediction (Papadopoulos et al., 2002b, ; Angelopoulos and Bates, , 2021; Straitouri et al., , 2023) Model agnostic Uncertainty Estimation for CATEs with weighted Conformal Prediction (Lei and Candès, , 2021; Tibshirani et al., , 2019) and conformal meta-learners (Alaa et al., , 2023) Uncertainty in Policy Value (Wang et al., , 2017) and Conformal Off-Policy Evaluation for Outcomes (Taufiq et al., , 2022)
Table 3: Overview of Methodological Approaches to Address Key Challenges of ADM Systems in the Public Sector in Risk Prediction, Counterfactual Prediction, and Off-Policy Learning.
Figure 3: Key Questions for Policy Makers in Selecting Risk Prediction as the Modeling Approach. Example inspired by algorithmic predictions of acute gastrointestinal bleeding (Alur et al., , 2023)
Figure 4: Key Questions for Policy Makers in Selecting Counterfactual Prediction as the Modeling Approach. Example inspired by child abuse hotline screening in Allegheny County (Coston et al., , 2020; Chouldechova et al., , 2018)
Figure 5: Key Questions for Policy Makers in Selecting Policy Learning as the Modeling Approach. Example inspired by geographical matching of refugees to improve integration outcomes (Bansak et al., , 2018)
Appendix B CATE Estimation Methods
@@
Alternative methods directly estimate the CATE function by first constructing pseudo-outcomes of the treatment effects (Künzel et al., , 2019; Caron et al., , 2022; Curth and van der Schaar, , 2021). One prominent variant is the X-learner (Künzel et al., , 2019), an extension of the T-learner, which can also be regarded as a special case of the RA-learner (Curth and van der Schaar, , 2021). The Doubly-Robust learner (DR-learner) employs pseudo-outcomes constructed from both the conditional outcomes $\hat{\mu}_{t}(x)$ and the propensity scores $\hat{\pi}(x)$ (Kennedy, 2020b, ). DR estimators have a long history in causal inference and missing data imputation (Kang and Schafer, , 2007; Bang and Robins, , 2005; Robins et al., , 1994; Funk et al., , 2011) and offer the advantage of consistency as long as either the propensity score model or conditional outcome models are correctly specified. Finally, the R-learner (Nie and Wager, , 2020; Foster and Syrgkanis, , 2023) involves the formulation a specific loss function after fitting several nuisance functions that can be separately minimized and regularized to estimate the CATE, drawing inspiration from the Robinson decomposition (Robinson, , 1988). There also exists CATE estimation methods that adapt specific ML models (Jacob, , 2021). For instance, causal forests, introduced by Athey et al., 2019b ; Wager and Athey, (2018), resemble the R-learner (Caron et al., , 2022; Post et al., , 2024; Oprescu et al., , 2019).
@@
Appendix C Off-Policy Learning